# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ 5.1 Toy examples: linear Gaussian low-level causal models
Figure 3: Comparison between learned and analytical $\bm{\tau}_{1}$ and ${\bm{\omega}}_{1}$ parameters. We show the average cosine similarity to the analytical solutions over 20 runs. Each run corresponds to one draw of adjacency matrix parameters. The shaded areas correspond to the range between the minimum and maximum values.
@@
We first test the proposed Algorithm 1 by sampling from a linear Gaussian low-level model, rather than a simulation. We construct linear models of the form shown in Proposition 3.3 by drawing the non-zero entries in the adjacency matrix uniformly from the interval $[-1,1]$. We learn a targeted causal reduction with two high-level variables: the target $Y$ and its single cause $Z$. Comprehensive experimental details can be found in App. F.
Fig. 3 compares the learned parameters $\bm{\tau}_{1}$ and ${\bm{\omega}}_{1}$ to the analytical solutions (7) and (8). We observe that, for these low-level models which meet the linear Gaussian assumption in Section 3, the learning algorithm converges to the global optimum.
Linear chain.
Figure 4: Comparison of learned parameters $\bm{\tau}_{1}$ and ${\bm{\omega}}_{1}$ for chain graphs with and without homogeneity regularization. The normalized $\bm{\tau}_{1}$ parameters ($(\bm{\tau}_{1})_{k}/\sum_{k^{\prime}}(\bm{\tau}_{1})_{k^{\prime}}$, where $k$ corresponds to the low-level variable index along the chain) are averaged over 20 runs. Each run corresponds to a draw of the adjacency parameters. Similarly, we depict the normalized ${\bm{\omega}}_{1}$ parameters. We compare two training settings: with and without homogeneity regularization (9) ($\eta_{h}=0$ or $\eta_{h}=1$, respectively). The inset shows the applied TCR scheme.
Figure 5: Double well experiment. Comparison of samples and inferred causal factors in the double well experiment. (a) Depicts the experimental setup with a ball moving in a double well potential subject to linear friction. (b) Displays the pushforward density of the high-level cause for the two settings: one where no intervention is applied (unintervened), and the other with an applied shift intervention. (c) and (d) represent the learned parameters, $\tau$ and $\omega$, respectively. (e) Illustrates the conditional mean $\mathbb{E}[Y\mid Z]$ for the inferred high-level causal mechanism. (f) and (g) show samples from the unintervened setting and the corresponding estimated density. (h) Depicts the estimated density for the intervened setting.
To illustrate the behavior of TCR under homogeneity regularization (9), we consider a linear chain causal graph. In a chain graph, we constrain the adjacency matrix such that all nodes are connected along a chain of edges, with the target variable being the leaf node ($X_{1}\to X_{2}\to\ldots\to Y$).
As depicted in Fig. 4, TCR attributes more significance to the low-level variables that are close ancestors in the causal graph (see interpretation in Sec. 3.3). By introducing homogeneity regularization (9), $\bm{\tau}_{1}$ and ${\bm{\omega}}_{1}$ become more congruent, enhancing interpretability. However, this regularization can reduce consistency (as elaborated in App. F).
5.2 ODE simulation: Double Well
@@
For a simulation based on an ODE system, we learn a targeted reduction of a ball moving in a double well potential under linear friction, as shown in Fig. 5. The state vector $\mathbf{X}$ encodes the $x$-position and velocity in $x$-direction of the ball at each time steps of the simulation. As shift-interventions, we apply small random shifts of the ball’s velocity at each simulation time step, mimicking an applied external force. Initially, the ball starts on the left-hand side of the potential and starts oscillating. Since the ball experiences friction, it ends up in either the left or right minimum of the potential. The friction is relatively strong, such that, depending on the initial conditions and applied shift interventions, the ball either stays in the left well or crosses the middle hump once and stays in the right well (see Fig. 5(f)). We learn a simple targeted reduction with a single cause $Z$ that explains the target $Y$. Further details about the ODE system and the learning parameters are given in App. F.
@@
The learned $\bm{\tau}_{1}$ and ${\bm{\omega}}_{1}$ parameters are shown in Fig. 5(c) and (d). Note that due to the negative slope of the learned high-level causal mechanism (Fig. 5(e)), negative $Z$-values are mapped to predicting the right well as the final position, and vice versa. The $\bm{\tau}_{1}$ and ${\bm{\omega}}_{1}$ parameters for velocity are such that the larger the velocity is to the right, the lower $Z$ and therefore the higher the predicted target $Y$. Similarly, for the position parameter: the more negative the position just before the critical point of the ball crossing the hump, the higher the probability of predicting to stay in the left well. This corresponds to the correct dynamics of the system and also identifies the main drivers that influence the outcome $Y$. Note that TCR does not focus on the part of $\mathbf{X}$ which best predicts the final state of the system—like the position just before the end of the simulation. It rather highlights the variables which have the most impact on the target when they are intervened on, emphasizing the decisive time span when the ball either crosses the middle hump or stays in the left well.
6 RELATED WORK
@@
Desiderata for CMR have been addressed theoretically by several works, in particular in the context of CFL (Chalupka et al., 2014, 2016), and subsequently with the notion of exact transformations (Rubenstein et al., 2017) and a more strongly constrained subclass: causal abstractions (Beckers and Halpern, 2019). An alternative framework for composing abstractions of finite models has been proposed by Rischel and Weichwald (2021). However, only few works have addressed how to build high-level representations from the low-level system data only. A line of works focuses on language models (Geiger and Straehle, 2020; Geiger et al., 2021, 2023a), where high-level variables and interpretations are readily available.
@@
We start from the opposite direction and develop a general approach to build the high-level abstraction from the ground-up. Such a construction is done in CFL (Chalupka et al., 2014, 2016), where high-dimensional microscopic variables are turned into discete high-level variables. Zennaro et al. (2023) addressed this question in the context of finite and discrete domains, by minimizing the maximum Jensen-Shannon divergence over a finite set of perfect intervention distributions. In contrast with most works on CMR, our framework is fully compatible with imperfect (soft intervention) at the low level, which are more realistic and interpretable perturbations of many real-world systems than hard interventions. Soft interventions have been used for language model alignment (Geiger et al., 2023b), and their theoretical compatibility with the abstraction framework has been investigated by Massidda et al. (2023). Our approach aims at approximating an exact transformation, and is thus a relaxation of this setting.
@@
Our approach also relates to the search for optimal interventional or counterfactual manipulations to steer the output of a system to a particular value or distribution (Amos et al., 2018; Besserve et al., 2020) or to best explain an observation (Budhathoki et al., 2022; Von Kügelgen et al., 2023). We are in a way also selecting particular manipulations, but through the choice of dimensionality reduction $\omega$, such that they are interpretable at a high level.
Finally, our approach relates to several works in causal representation learning, which have addressed identifiability of latent causal models from observational data, with (Liang et al., 2023) or without (Squires et al., 2023; von Kügelgen et al., 2023) assumptions on the latent causal graph. In contrast to those works, TCR does not assume an injective mapping of the mapping of the observations to the latent variables, such that the high-level model typically losses information relative to the low-level model.
7 DISCUSSION
Summary.
We introduce a novel approach for understanding complex simulations by learning high-level causal explanations from low-level models. We propose the Targeted Causal Reduction (TCR) framework, which leverages interventions in the low-level model to obtain simplified, high-level representations of the causes of a target phenomenon. We formulate the intervention-based consistency constraint as an information theoretic learning objective, and provide theoretical interpretations and justifications for it. By further enforcing linearity and Gaussianity assumptions, and using shift interventions, we provide analytical solutions and study their uniqueness. This provides insights into TCR’s governing principles and potential weaknesses, which we address through regularization. We provide an algorithm for linear TCR and show it can effectively uncover the key causal factors influencing a phenomenon of interest. We demonstrate the algorithm’s capabilities through applications to both synthetic and scientific models, highlighting its potential for addressing the challenges posed by increasingly complex systems in scientific research.
Limitations and future work.
To foster interpretability and tractability, we made Gaussian approximations and used linear $\bm{\tau}$ and ${\bm{\omega}}$ maps. While this has computational and interpretability benefits, this may be too limiting for some complex simulations, and future work should explore more flexible approaches. Additionally, our method relies on performing a large number of interventions in simulation runs, which represents an additional cost in the context of large-scale simulation. How to make the algorithm scale to this setting is left to future work.
Supplementary Materials
Appendix A SUPPLEMENTAL BACKGROUND
A.1 Numerical schemes for simulations
Methods for the numerical approximations of scientific models is a broad area spanning multiple fields. We provide here a few elements based on a 1D example to justify how these models relate to SCMs. The Euler method (Biswas et al., 2013), can be used to approximate a 1D ODE of the form
@@ A.1 Numerical schemes for simulations
This approach generalizes to explicit numerical schemes for multivariate ODE and SDEs where the state variable $\mathbf{X}$ lives in $\mathbb{R}^{n}$. As an illustration we can take the following class of SDE models
@@
with $\mathbf{F}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, $\bm{\sigma}_{W}=\mathbb{R}^{(n\times n)}$, and $\mathbf{W}$ a $n$-dimentional standard Brownian motion. This leads to the scheme
@@
where the $\mathbf{U}_{k}$ are now multivariate Gaussian variables, whose components may or may not be independent depending on the choice of the matrix $\bm{\sigma}_{W}$. If the exogenous components are independent, the variables can be described by a standard SCM as introduced in main text. If the exogenous components are dependent, the variables can be described by a more general notion of SCM, allowing hidden confounding (Bongers et al., 2021).
@@
A.2 Reduction of the Euler scheme for a system of point masses
@@
We can approximate each equation to estimate interatively the $x$ and $y$ components of the speed of individual point masses in the system, using a small time-step $\Delta t$, such that we get the discrete time estimates $\hat{v}_{x,k}[n]\approx v_{x,k}(n\Delta t)$ and $\hat{v}_{y,k}[n]\approx v_{y,k}(n\Delta t)$ satisfying
@@
To make the notation compatible with those use used in our TCR framework, we can gather all speed variables in a high-dimensional vector $\mathbf{X}$ and all external force variables in a vector $\mathbf{i}$, the high-level causal model is thus generated by a linear $\tau$-map and linear $\omega$ map for shift interventions, taking the form of the exact transformation
@@ A.3 Constructive transformations
Appendix B PROOF OF MAIN TEXT RESULTS
B.1 Proof of Proposition 3.1
@@ Appendix B PROOF OF MAIN TEXT RESULTS
$\mathcal{L}_{cons}\!=\!\mathbb{E}_{i\sim P(\mathbf{i})}\!\!\left[KL\left(\widehat{P}_{\tau}^{(\mathbf{i})}\left(\mathbf{Z}\right)||{P}^{(\omega(\mathbf{i}))}\left(\mathbf{Z}\right)\right)+\mathbb{E}_{\mathbf{z}\sim\widehat{P}_{\tau}^{(\mathbf{i})}\left(\mathbf{Z}\right)}\!\!\left[KL\left(\widehat{P}_{\tau}^{(\mathbf{i})}\left({Y}|\mathbf{z})\right)||{P}^{(0)}\left({Y}|\mathbf{z}\right)\right)\right]\right]\,,$ (10)
@@
$\mathcal{L}_{rel}\!=\!\mathbb{E}_{\mathbf{i}\sim P(\mathbf{i})}\!\left[KL\left(\widehat{P}^{(\mathbf{i})}\left(Y\right)||{P}^{(\omega(\mathbf{i}))}\left(Y\right)\right)\right]\!\leq\!\mathcal{L}_{cons}\,.$ (11)
@@
Invariance to reparameterizations.
We assume a reparametrization $\rho$ designed according to the framework introduced in Appendix D.1. By invariance of the KL divergence to invertible transformations, we have equality between the KL associated to the two different reductions $(\tau,\omega)$ and $(\rho\circ\tau,\psi\circ\omega)$:
$KL\left(\widehat{P}_{\tau}^{(\mathbf{i})}({Y},{\mathbf{Z}})||P_{\mathcal{H},\bm{\gamma}}^{(\omega(\mathbf{i}))}(Y,\mathbf{Z})\right)=KL\left(\tilde{\rho}_{\#}[\widehat{P}_{\tau}^{(\mathbf{i})}({Y},{\mathbf{Z}})]||\tilde{\rho}_{\#}[P_{\mathcal{H},\bm{\gamma}}^{(\omega(\mathbf{i}))}(Y,\mathbf{Z})]\right)=KL\left(\widehat{P}_{\rho\circ\tau}^{(\mathbf{i})}({Y},{\mathbf{Z}})||P_{\mathcal{H},\bm{\gamma}^{\prime}}^{(\psi\circ\omega(\mathbf{i}))}(Y,\mathbf{Z})\right)\,.$
The transformation $(\rho,\psi)$ thus leaves $\mathcal{L}_{cons}$ invariant.
@@
$KL$ $\left(\widehat{P}^{(\mathbf{i})}({Y},{\mathbf{Z}})||P^{(\omega(\mathbf{i}))}({Y},{\mathbf{Z}})\right)$
@@
$=KL_{Z}\left(\widehat{P}^{(\mathbf{i})}\left(\mathbf{Z}\right)||{P}^{(\omega(\mathbf{i}))}\left(\mathbf{Z}\right)\right)+\mathbb{E}_{z\sim\widehat{P}^{(\mathbf{i})}\left(\mathbf{Z}\right)}\left[KL_{Y}\left(\widehat{P}^{(\mathbf{i})}\left({Y}|{\mathbf{Z}}=z\right)||{P}^{(0)}\left({Y}|{\mathbf{Z}}=z\right)\right)\right]$
$=KL_{Z}\left(\widehat{P}^{(\mathbf{i})}\left({\mathbf{Z}}\right)||{P}^{(\omega(\mathbf{i}))}\left(\mathbf{Z}\right)\right)+KL_{Y,Z}\left(\widehat{P}^{(\mathbf{i})}\left(\widehat{Y},{\mathbf{Z}}\right)||{P}^{(0)}\left({Y}|{\mathbf{Z}}\right)\widehat{P}^{(\mathbf{i})}\left(\mathbf{Z}\right)\right)\,.$
The first term is a cause consistency loss (same principle but for that variable only). The second term can be thought of as a mechanism consistency, where we use the ground truth low-level cause distribution to probe the similarity of the outputs of the “true” (in fact, the conditional distribution) and approximate mechanism. Our interpretability choice prevents the high-level mechanism to be intervened on, so a single stochastic map (i.e. a Markov kernel) must fit at best all the sampled experimental conditionals.
@@
$\mathcal{L}_{rel}=\mathbb{E}_{\mathbf{i}\sim p(\mathbf{i})}\left[KL_{Y}\left(\widehat{P}^{(\mathbf{i})}\left(Y\right)||{P}^{(\omega(\mathbf{i}))}\left(Y\right)\right)\right]\,.$
Note: In the Gaussian 1D case, this gives
$\frac{1}{2}\mathbb{E}_{\mathbf{i}\sim p(\mathbf{i})}\left[\frac{({\mu}_{Y}+\alpha{\bm{\omega}}^{\top}\mathbf{i}-\widehat{\mu}^{(\mathbf{i})}_{Y})^{2}}{\sigma_{Y}^{2}}+\frac{{\widehat{\sigma^{2}}}^{(\mathbf{i})}_{Y}}{\sigma_{Y}^{2}}-\ln{\left(\frac{{\widehat{\sigma^{2}}}^{(\mathbf{i})}_{Y}}{\sigma_{Y}^{2}}\right)}-1\right]\,.$
@@
$KL$ $\left(\widehat{P}^{(\mathbf{i})}({Y},{\mathbf{Z}})||P^{(\omega(\mathbf{i}))}({Y},{\mathbf{Z}})\right)$
$=KL_{Y}\left(\widehat{P}^{(\mathbf{i})}\left(Y\right)||{P}^{(\omega(\mathbf{i}))}\left(Y\right)\right)+\mathbb{E}_{y\sim\widehat{P}^{(\mathbf{i})}\left(Y\right)}\left[KL_{Y}\left(\widehat{P}^{(\mathbf{i})}\left({\mathbf{Z}}|{Y}=y\right)||{P}^{(\omega(\mathbf{i}))}\left({\mathbf{Z}}|{Y}=y\right)\right)\right]$

### v2

@@ 5.1 Toy examples: linear Gaussian low-level causal models
Figure 4: Spring-mass system experiment. (a) Simulated system of four point masses with different weights connected by springs and with random initial velocity (blue arrows). The target of the simulation is the center of mass speed in $(1,1)$-direction. (b-e) Learned $\bm{\tau}$- and ${\bm{\omega}}$-weights corresponding to velocity components in $x$- and $y$-direction for a TCR with two high-level variables. The learned high-level mechanism is $f(\mathbf{Z})\approx-0.226Z_{1}+0.220Z_{2}$. (f) Comparison between masses and learned omega weights. For the first high-level variable the mean omega weights corresponding to the $x$-direction are shown and for the second variable those for the $y$-direction. (g-j) Example trajectory for an unintervened system.
@@
We first test TCR by sampling from a linear Gaussian low-level model, rather than a simulation. We construct linear models of the form shown in Prop. 3.4 by drawing the non-zero entries in the adjacency matrix uniformly from the interval $[-1,1]$. We learn a targeted causal reduction with two high-level variables: the target $Y$ and its single cause $Z$. Fig. 2(a) compares the learned $\bm{\tau}_{1}$ and ${\bm{\omega}}_{1}$ to the analytical solutions (9) and (10). We observe that, for these low-level models meeting the linear Gaussian assumption in Section 3, the learning algorithm converges to the global optimum.
Two-branch model.
To investigate the behavior of TCR with multiple high-level variables, we consider a low-level model with two branch causal structure (Fig. 2(b)). With regularization for overlap (12) and balancing (13), the learned high-level variables correspond to the two branches. Within each branch, the reduction behaves as described for the linear chain in Ex. 3.5, where $\tau$ focusses on the direct parent of the target and $\omega$ is spread across all variables in the chain. Comprehensive experimental details are given in App. F.
5.2 Double Well
@@
For a simulation based on an ODE system, we learn a targeted reduction of a ball moving in a double well potential under linear friction, as shown in Fig. 3. The state vector $\mathbf{X}$ encodes the $x$-position and velocity in $x$-direction of the ball at each time steps of the simulation. As shift-interventions, we apply small random shifts of the ball’s velocity at each simulation time step, mimicking an applied external force. Initially, the ball starts on the left-hand side of the potential and starts oscillating. Since the ball experiences friction, it ends up in either the left or right minimum of the potential. The friction is relatively strong, such that, depending on the initial conditions and applied shift interventions, the ball either stays in the left well or crosses the middle hump once and stays in the right well (see Fig. 3(f)). We learn a simple TCR with a single cause $Z$ that explains the target $Y$. Further details about the nonlinear ODE system and training are given in App. F.2.
@@
The learned TCR parameters are shown in Fig. 3(c, d). The $\bm{\tau}_{1}$ and ${\bm{\omega}}_{1}$ parameters for velocity are such that the larger the velocity is to the right, the higher $Z$ and therefore the higher the predicted target $Y$, where positive $Y$ correspond to the right well and negative to the left. Similarly, for the position parameter: the more negative the position just before the critical point of the ball crossing the hump, the higher the probability of predicting to stay in the left well. This corresponds to the correct dynamics of the system and also identifies the main drivers that influence the outcome $Y$. Fig. 3(e) shows how the learned TCR separates the phase space into simulations with enough momentum to the right to make it over the hump (pink) and those without (turquoise).
Note that TCR does not focus on the part of $\mathbf{X}$ which best predicts the final state of the system—like the position just before the end of the simulation. It rather highlights the variables which have the most impact on the target when they are intervened on, emphasizing the decisive time span when the ball either crosses the middle hump or stays in the left well.
5.3 Spring-Mass System
We simulate a two-dimensional system of four point masses with different weights connected by springs to their respective nearest neighbors, similar to the motivating example introduced in Sec. 1. Initially, the masses are arranged in a rectangle in space such that the springs are at rest length. The masses have a random initial velocity, as shown in Fig. 4(a). As interventions, we apply random shifts to the velocities in $x$- and $y$-direction of each mass. The target of the simulation is the center of mass speed in $(1,1)$-direction. We learn a TCR with two high-level causes. The full experimental details are given in App. F.3.
While the velocities of the individual masses are coupled, the center of mass velocities in $x$- and $y$-direction of the system as a whole are independent, since the system is freely moving in space. The learned TCR correctly identifies these as the two independent causes of the target, with variable $Z_{1}$ corresponding to the $y$-direction and $Z_{2}$ to the $x$-direction. On average, each mass receives a similar shift in velocity through the applied interventions. However, since the masses are different, the shifts correspond to different contributions to the momentum of the system as a whole impacting the target. This is reflected in the relative weights of the learned maps being proportional to the weight of each point mass, as shown in Fig 4(f).
A second experimental setting with two groups of interconnected masses is shown in App. F.4.2, demonstrating a TCR learning independent causes along the mass index.
6 DISCUSSION
We introduce a novel approach for understanding complex simulations by learning high-level causal explanations from low-level models. Our Targeted Causal Reduction (TCR) framework leverages interventions to obtain simplified, high-level representations of the causes of a target phenomenon. We formulate the intervention-based consistency constraint as an information theoretic learning objective, which favors the most causally relevant explanations of the target. Under linearity and Gaussianity assumptions, we provide analytical solutions and study their uniqueness, which provides insights into TCR’s governing principles. One key assumption to obtain identifiability is that the leaf node, the target, is observed. However, this is to the best of our knowledge the first identifiability proof for a general class of CMR for which the high-level variables are continuous and partially unknown. Notably, the $n$-cause TCR provides a form of causal independent component analysis akin to the work of Wendong et al. (2023) but in a non-invertible setting and with a one dimensional target. We provide an algorithm for linear TCR and show it can effectively uncover the key causal factors influencing a phenomenon of interest. We demonstrate TCR on both synthetic models and scientific simulations, highlighting its potential for addressing the challenges posed by increasingly complex systems in scientific research.
While we develop a CMR framework to learn high-level explanations for simulations, the simulation itself does not have to be explicitly formulated as a causal model and the causal relationships between variables in $\mathbf{X}$ do not have to be known a priori. The only additional element needed to learn TCR is a notion of shift-interventions. We think that most scientific simulations based on differential equations naturally allow for a reasonable notion of shift interventions.
Limitations and future work.
To foster interpretability and tractability, we made Gaussian approximations and used linear $\bm{\tau}$ and ${\bm{\omega}}$ maps. While this has clear benefits, this may be too limiting for some complex simulations, and future work should explore more flexible approaches. Additionally, our method relies on performing a large number of interventions in simulation runs, which represents an additional cost in the context of large-scale simulation. How to make the algorithm scale to this setting is left to future work.
Acknowledgements.
We thank Sergio Hernan Garrido Mejia and Yuchen Zhu for insightful discussions. This publication was supported by the German Federal Ministry of Education and Research (BMBF) through the Tübingen AI Center (FKZ 01IS18039A) and by the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) through the Machine Learning Cluster of Excellence (EXC 2064/1, project 390727645).
Appendix A SUPPLEMENTAL RELATED WORK
@@
Desiderata for CMR have been addressed theoretically by several works, in particular in the context of CFL (Chalupka et al., 2015, 2016), and subsequently with the notion of exact transformations (Rubenstein et al., 2017) and a more strongly constrained subclass: causal abstractions (Beckers and Halpern, 2019). An alternative framework for composing abstractions of finite models has been proposed by Rischel and Weichwald (2021). However, only few works have addressed how to build high-level representations from the low-level system data only. A line of works focuses on language models (Geiger et al., 2021, 2023), where high-level variables and interpretations are readily available.
@@
We start from the opposite direction and develop a general approach to build the high-level abstraction from the ground-up. Such a construction is done in CFL (Chalupka et al., 2015, 2016), where high-dimensional microscopic variables are turned into discrete high-level variables. Zennaro et al. (2023) addressed this question in the context of finite and discrete domains, by minimizing the maximum Jensen-Shannon divergence over a finite set of perfect intervention distributions. In contrast with most works on CMR, our framework is fully compatible with imperfect (soft intervention) at the low level, which are more realistic and interpretable perturbations of many real-world systems than hard interventions. Soft interventions have been used for language model alignment (Geiger et al., 2024), and their theoretical compatibility with the abstraction framework has been investigated by Massidda et al. (2023). Our approach aims at approximating an exact transformation, and is thus a relaxation of this setting.
@@
Our approach also relates to the search for optimal interventional or counterfactual manipulations to steer the output of a system to a particular value or distribution (Amos et al., 2018; Besserve et al., 2020) or to best explain an observation (Budhathoki et al., 2022; von Kügelgen et al., 2023b). We are in a way also selecting particular manipulations, but through the choice of dimensionality reduction $\omega$, such that they are interpretable at a high level.
Finally, our approach relates to several works in causal representation learning, which have addressed identifiability of latent causal models from observational data, with (Wendong et al., 2023) or without (Squires et al., 2023; von Kügelgen et al., 2023a) assumptions on the latent causal graph. In contrast to those works, TCR does not assume an injective mapping of the mapping of the observations to the latent variables, such that the high-level model typically losses information relative to the low-level model.
Appendix B SUPPLEMENTAL BACKGROUND
B.1 Numerical schemes for simulations
Methods for the numerical approximations of scientific models is a broad area spanning multiple fields. We provide here a few elements based on a 1D example to justify how these models relate to SCMs. The Euler method (Euler, 1794), can be used to approximate a 1D ODE of the form
@@ B.1 Numerical schemes for simulations
This approach generalizes to explicit numerical schemes for multivariate ODE and SDEs where the state variable $\mathbf{X}$ is an element of $\mathbb{R}^{n}$. As an illustration, we can take the following class of SDE models
@@
with $\mathbf{F}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, $\bm{\sigma}_{W}=\mathbb{R}^{(n\times n)}$, and $\mathbf{W}$ a $n$-dimensional standard Brownian motion. This leads to the scheme
@@
where the $\mathbf{U}_{k}$ are now multivariate Gaussian variables, whose components may or may not be independent depending on the choice of the matrix $\bm{\sigma}_{W}$. If the exogenous components are independent, the variables can be described by a standard SCM as introduced in the main text. If the exogenous components are dependent, the variables can be described by a more general notion of SCM, allowing hidden confounding (Bongers et al., 2021).
@@
B.2 Reduction of the Euler scheme for a system of point masses
@@
We can approximate each equation to iteratively estimate the $x$ and $y$ components of the speed of individual point masses in the system, using a small time-step $\Delta t$, such that we get the discrete time estimates $\hat{v}_{x,k}[n]\approx v_{x,k}(n\Delta t)$ and $\hat{v}_{y,k}[n]\approx v_{y,k}(n\Delta t)$ satisfying
@@
To make the notation compatible with that used in our TCR framework, we can gather all speed variables in a high-dimensional vector $\mathbf{X}$ and all external force variables in a vector $\mathbf{i}$, the high-level causal model is thus generated by a linear $\tau$-map and linear $\omega$ map for shift interventions, taking the form of the exact transformation
@@ B.3 Constructive transformations
Appendix C PROOF OF MAIN TEXT RESULTS
C.1 Proof of Proposition 3.1
@@ Appendix C PROOF OF MAIN TEXT RESULTS
$\mathcal{L}_{\mathrm{cons}}\!=\!\mathbb{E}_{i\sim P(\mathbf{i})}\!\!\left[\mathrm{KL}\left(\widehat{P}_{\tau}^{(\mathbf{i})}\left(\mathbf{Z}\right)||{P}^{(\omega(\mathbf{i}))}\left(\mathbf{Z}\right)\right)+\mathbb{E}_{\mathbf{z}\sim\widehat{P}_{\tau}^{(\mathbf{i})}\left(\mathbf{Z}\right)}\!\!\left[\mathrm{KL}\left(\widehat{P}_{\tau}^{(\mathbf{i})}\left({Y}|\mathbf{z})\right)||{P}^{(0)}\left({Y}|\mathbf{z}\right)\right)\right]\right]\,,$ (15)
@@
$\mathcal{L}_{\mathrm{rel}}\!=\!\mathbb{E}_{\mathbf{i}\sim P(\mathbf{i})}\!\left[\mathrm{KL}\left(\widehat{P}^{(\mathbf{i})}\left(Y\right)||{P}^{(\omega(\mathbf{i}))}\left(Y\right)\right)\right]\!\leq\!\mathcal{L}_{\mathrm{cons}}\,.$ (16)
@@
Invariance to reparameterizations. We assume a reparametrization $\rho$ designed according to the framework introduced in Appendix D.1. By invariance of the KL divergence to invertible transformations, we have equality between the KL associated to the two different reductions $(\tau,\omega)$ and $(\rho\circ\tau,\psi\circ\omega)$:
$\mathrm{KL}\left(\widehat{P}_{\tau}^{(\mathbf{i})}({Y},{\mathbf{Z}})||P_{\mathcal{H},\bm{\gamma}}^{(\omega(\mathbf{i}))}(Y,\mathbf{Z})\right)=\mathrm{KL}\left(\tilde{\rho}_{\#}[\widehat{P}_{\tau}^{(\mathbf{i})}({Y},{\mathbf{Z}})]||\tilde{\rho}_{\#}[P_{\mathcal{H},\bm{\gamma}}^{(\omega(\mathbf{i}))}(Y,\mathbf{Z})]\right)=\mathrm{KL}\left(\widehat{P}_{\rho\circ\tau}^{(\mathbf{i})}({Y},{\mathbf{Z}})||P_{\mathcal{H},\bm{\gamma}^{\prime}}^{(\psi\circ\omega(\mathbf{i}))}(Y,\mathbf{Z})\right)\,.$
The transformation $(\rho,\psi)$ thus leaves $\mathcal{L}_{\mathrm{cons}}$ invariant.
@@
$\mathrm{KL}$ $\left(\widehat{P}^{(\mathbf{i})}({Y},{\mathbf{Z}})||P^{(\omega(\mathbf{i}))}({Y},{\mathbf{Z}})\right)$
@@
$=\mathrm{KL}_{Z}\left(\widehat{P}^{(\mathbf{i})}\left(\mathbf{Z}\right)||{P}^{(\omega(\mathbf{i}))}\left(\mathbf{Z}\right)\right)+\mathbb{E}_{z\sim\widehat{P}^{(\mathbf{i})}\left(\mathbf{Z}\right)}\left[\mathrm{KL}_{Y}\left(\widehat{P}^{(\mathbf{i})}\left({Y}|{\mathbf{Z}}=z\right)||{P}^{(0)}\left({Y}|{\mathbf{Z}}=z\right)\right)\right]$
$=\mathrm{KL}_{Z}\left(\widehat{P}^{(\mathbf{i})}\left({\mathbf{Z}}\right)||{P}^{(\omega(\mathbf{i}))}\left(\mathbf{Z}\right)\right)+\mathrm{KL}_{Y,Z}\left(\widehat{P}^{(\mathbf{i})}\left(\widehat{Y},{\mathbf{Z}}\right)||{P}^{(0)}\left({Y}|{\mathbf{Z}}\right)\widehat{P}^{(\mathbf{i})}\left(\mathbf{Z}\right)\right)\,.$
We call the first term cause consistency loss, as it matches the definition of a consistency loss but for cause variables only. The second term can be thought of as a mechanism consistency loss, where we use the ground truth low-level cause distribution to probe the similarity of the outputs of the “true” (in fact, the conditional distribution) and approximate mechanism. Our interpretability choice prevents the high-level mechanism from being intervened on, so a single stochastic map (i.e. a Markov kernel) must fit at best all the sampled experimental conditionals.
@@
$\mathcal{L}_{\mathrm{rel}}=\mathbb{E}_{\mathbf{i}\sim p(\mathbf{i})}\left[\mathrm{KL}_{Y}\left(\widehat{P}^{(\mathbf{i})}\left(Y\right)||{P}^{(\omega(\mathbf{i}))}\left(Y\right)\right)\right]\,.$
Note: In the Gaussian 1D case, the formula for the causal relevance loss is
$\mathcal{L}_{\mathrm{rel}}=\frac{1}{2}\mathbb{E}_{\mathbf{i}\sim p(\mathbf{i})}\left[\frac{({\mu}_{Y}+\alpha{\bm{\omega}}^{\top}\mathbf{i}-\widehat{\mu}^{(\mathbf{i})}_{Y})^{2}}{\sigma_{Y}^{2}}+\frac{{\widehat{\sigma^{2}}}^{(\mathbf{i})}_{Y}}{\sigma_{Y}^{2}}-\ln{\left(\frac{{\widehat{\sigma^{2}}}^{(\mathbf{i})}_{Y}}{\sigma_{Y}^{2}}\right)}-1\right]\,.$
@@
$\mathrm{KL}$ $\left(\widehat{P}^{(\mathbf{i})}({Y},{\mathbf{Z}})||P^{(\omega(\mathbf{i}))}({Y},{\mathbf{Z}})\right)$
$=\mathrm{KL}_{Y}\left(\widehat{P}^{(\mathbf{i})}\left(Y\right)||{P}^{(\omega(\mathbf{i}))}\left(Y\right)\right)+\mathbb{E}_{y\sim\widehat{P}^{(\mathbf{i})}\left(Y\right)}\left[\mathrm{KL}_{Y}\left(\widehat{P}^{(\mathbf{i})}\left({\mathbf{Z}}|{Y}=y\right)||{P}^{(\omega(\mathbf{i}))}\left({\mathbf{Z}}|{Y}=y\right)\right)\right]$